# AI-based particle track identification in scintillating fibres read out with imaging sensors

### v1

@@ Appendix A Additional results
This appendix provides supplementary results, demonstrating the impact of varying thresholds on Binary Cross-Entropy (BCE) and Kullback-Leibler Divergence (KLD) loss values of the variational autoencoder (VAE). The results highlight how using these loss functions individually or in combination yields different sets of anomalies. Specifically, the following sections focus on anomaly selection based on the criteria detailed below:
@@ A.3 Anomaly on both BCE and KLD
Figure 12 shows the total sum of frames selected and not selected as anomaly candidates based on different thresholds applied to either the BCE or KLD loss values, while Table 5 presents the total number of frames selected in each case, categorized by the number of photon counts in the frame. The results suggest that combining both BCE and KLD performs similarly to using BCE alone. Future work could explore more sophisticated methods of combining these loss functions to leverage their strengths and improve anomaly detection efficiency.
@@ Acknowledgments
We would like to thank the AQUA-LAB at EPFL for their help and fruitful discussions about the data acquisition for this project. Further we would like to thank them for the provision of the SwissSPAD2 imaging sensor.

### v3

@@ Appendix A Additional results
This appendix provides supplementary results, demonstrating the impact of varying thresholds on BCE and KLD loss values of the VAE. The results highlight how using these loss functions individually or in combination yields different sets of anomalies. Specifically, the following sections focus on anomaly selection based on the criteria detailed below:
@@ A.3 Anomaly on both BCE and KLD
Figure 16 shows the total sum of frames selected and not selected as anomaly candidates based on different thresholds applied to either the BCE or KLD loss values, while Table 5 presents the total number of frames selected in each case, categorised by the number of photon counts in the frame. The results suggest that combining both BCE and KLD performs similarly to using BCE alone. Future work could explore more sophisticated methods of combining these loss functions to leverage their strengths and improve anomaly detection efficiency.
@@ Acknowledgments
We would like to thank the Edoardo Charbon and Claudio Bruschini and the Advanced Quantum Architecture Laboratory at EPFL for their help and fruitful discussions about the data acquisition for this project. Further, we would like to thank them for the provision of the SwissSPAD2 imaging sensor and the dataset used for testing the developed analysis method.